Engine oil supplied through a 2cm diameter thin walled copper pipe. k-385w/kg,c is to be heated from 20C to 60C at a rate of 0.3kg/s in a double-pipe heat exchanger by condensing steam at a temperature of 130C, which is supplied through an outer pope with a diameter of 2 cm. if the heat exchager is well insulated determine the length of the heat exchanger
Answer #1

Tsat =130°c 02=130-60 Foc Oo-Tsat-Ti To=60c =130-20 Mail = 0.349/ -lloc Ti = 20°C - 2cm pipe 01-02 E 110- 70 (LMTD) of HeatRez PUD M Re PVDd xD m=8AV ex ID PvX/A) Re ambil la XD Property of oil ( Engine oil) fval Value taken from Property table -0.A thumbs up, pls !!